OpenStudy (lgbasallote):

when disproving universal quantifiers, a counterexample is stated. When proving existential quantifier, one example that suits the expression is all that's needed. What about in proving universal quantifiers?

OpenStudy (zugzwang):

\[\forall x f(x)\] Where f(x) is a truth function (could either be true or false) To prove it, get an arbitrary x, and show that f(x) is true.

OpenStudy (lgbasallote):

but even if you get arbitrary x, it does not mean it is applicable for all x

OpenStudy (zugzwang):

Yes it does...

OpenStudy (lgbasallote):

for axample what's given is \[Q(x) = x < 6\]

OpenStudy (anonymous):

When, proving a universal quantifier A is true for all B, you could disprove the converse statement A is untrue for some B.

OpenStudy (lgbasallote):

even if you say \[\text{At} \; Q(4): \\ 4 < 6 \\ \therefore \forall x Q(x) \equiv T\]

OpenStudy (lgbasallote):

this cannot be said...

OpenStudy (lgbasallote):

because at Q(7) 7 < 4 \(\forall\)x Q(x) \(\equiv\) F

OpenStudy (zugzwang):

The statement is false. You can produce a counterexample.

OpenStudy (lgbasallote):

that's what i'm sayin

OpenStudy (lgbasallote):

saying*

OpenStudy (zugzwang):

And you did not take an arbitrary value for x, you took a specific value.

OpenStudy (lgbasallote):

by arbitrary you mean..?

OpenStudy (zugzwang):

I think it best to illustrate via an example: To prove that For all integers, if it is odd, then its square is odd. \[\forall x \in Z, x \ is \ odd \rightarrow x^{2} \ is \ odd\] To prove it Let x be an arbitrary integer (THIS IS WHAT I MEAN) Suppose x is odd. Then x = 2k + 1 for some integer k. Then x^2 = (2k + 1)^2 = 4k^2 + 4k + 1 = 2(2k^2 + 2k) + 1 But 2k^2 + 2k is an integer, therefore x^2 is odd. QED

OpenStudy (lgbasallote):

that's already proving....

OpenStudy (lgbasallote):

i mean "proving" proving

OpenStudy (zugzwang):

I know, and what did, that's what I mean by taking an arbitrary value, you're not supposed to assign a value to it, as you did by assigning x = 4 and x = 7

OpenStudy (lgbasallote):

but what i did is what i meant

OpenStudy (lgbasallote):

that was what i was asking

OpenStudy (zugzwang):

I thought you were asking how to prove universal quantifiers...

OpenStudy (lgbasallote):

i was asking how to prove universal quantifiers using *numbers*

OpenStudy (zugzwang):

Well, that, I can't say for sure, unless you can show me an example of what to prove...

OpenStudy (lgbasallote):

Let P(x) = "x+ 1 > x". Find the truth value of \(\forall\)x P(x) if the domain consists of all real numbers

OpenStudy (zugzwang):

Hmm... I'm afraid I can only prove this using the method shown above... As you said, examples ain't gonna do it for proving universal quantifiers...

OpenStudy (lgbasallote):

i supose there's no other way then

OpenStudy (zugzwang):

Let x be a real number 1 > 0 Since x is a real number, by the addition property of inequality, x + 1 > 0 + x x + 1 > x It's possible to prove 1 > 0, but I take it your instructor is not quite that sadistic...

OpenStudy (lgbasallote):

instructor? i have no instructor

OpenStudy (zugzwang):

Then replace "your instructor is not quite that sadistic..." with "you're not quite that masochistic"